# Electron Magneto-Hydrodynamics in Graphene

### v1

@@ III.1 Boundary layer flow in the perturbative limit $\omega\tau\ll 1$
The no-penetration condition ensures that no fluid can cross the boundary whereas the no-slip condition ensures that the fluid has zero tangential velocity at the boundary surface. Furthermore, the stream function must decay to the bulk solution (4) far from the boundary. We look for a solution as an expansion in powers of $\omega\tau\ll 1$, i.e., $\psi=\sum_{n=0}^{\infty}(\omega\tau)^{n}\psi^{(n)}$. First we solve for the zero-order solution, corresponding to flow in the absence of a magnetic field. Solving the Stokes equation for $\omega\tau=0$ using the boundary conditions leads to the Poiseuille solution
@@ III.2 Exact numerical solution for the boundary layer flow
Figure 1: Streamlines in a semi-infinite channel with a no-slip boundary at $y=0$. The parameters are chosen as $gD=1.16$, $\omega\tau=0.3,1.2,4.5$. Solid lines correspond to the numerically exact solution described in Sec. III.2 and dashed lines correspond to the solution given by Eq. (4) in the infinite device.
@@
where $f_{m}$ and $A_{m,a}$ are complex coefficients. The index $a=1,2,3...$ enumerates “eigenvalues” $\lambda_{a}$, which can be real or complex. The real parts of $\lambda_{a}$ are positive by design of Eq. (14). In principle the index $a$ can be arbitrarily large, but due to the truncation it is never larger than $\sim 10$. Since the “Bernoulli term” is neglected, the Stokes equation (II) is a linear equation. Let us consider the Poiseuille solution without magnetic field,
@@
The boundary layer solution considered in this Section is valid for the no-slip boundary condition. However, the method can be easily extended to the case of an arbitrary slip length, considering that only Eqs. (III.2) depend on the slip length.
The boundary layer solution is extended to finite width channels in App. A. After the method is developed it is instructive to present plots for specific cases. For the numerical solution we set $gD=1.16$, which corresponds to the device described at the end of Sec. II. We truncate beyond $N_{m}=5$, which provides more than sufficient accuracy for the values of $\omega\tau$ we present. Plots of the streamlines obtained for $\omega\tau=0.3$, $1.2$ and $4.5$ are presented in Fig. 1. These values of $\omega\tau$ were selected as $\omega\tau=0.3$ has only real eigenvalues $\lambda_{a}$ and corresponds to the perturbative regime, $\omega\tau=1.2$ has a complex conjugate pair of eigenvalues $\lambda_{a}$ and corresponds to the device described in Eq. (II) and $\omega\tau=4.5$ has two complex conjugate pairs of eigenvalues $\lambda_{a}$. As the magnetic field increases ($\omega\tau$ increases) both the magnitude of the boundary effects and the width of the boundary layer increase, but in the cases considered the effective width of the boundary layer does not extend beyond $y/D\approx 4$. Thus, considering symmetry, it is appropriate to use the bulk solution in devices with width $W\gg 8D$.
@@ IV Numerical estimates, anti-Poiseuille flow
Following the numerical calculation of the stream function, we may calculate the the viscous dissipation rate and the Ohmic dissipation rate given by Eqs. (II). The most important parameter is the ratio $\dot{E}_{\nu}/\dot{E}_{\Omega}$, which is dependent on the device considered, so to provide calculated values we take the device with parameters presented in Eq. (II). This corresponds to $D=0.34\>\mu\text{m}$. We also take the lattice spacing $a=1.9\>\mu\text{m}$ as discussed in the end of Section II, resulting in $gD=1.16$. The ratio $\dot{E}_{\nu}/\dot{E}_{\Omega}$ is plotted in Fig. 2 versus the channel width $W$, for two values of the magnetic field amplitude, $B_{0}=75$ mT,$150$ mT, corresponding to $\omega\tau=1.2$, $2.4$ respectively. In these plots we assume that the viscosity is independent of the magnetic field, $\nu(B)=\nu(0)$. Asymptotic values of the ratio $\dot{E}_{\nu}/\dot{E}_{\Omega}$ at $W\to\infty$ are $0.16$ ($B_{0}=75$ mT) and $0.48$ ($B_{0}=150$ mT).
@@
Figure 2: Ohmic (blue line) and viscous (red line) dissipation as a function of the width of the channel normalized to the $\dot{E}_{0}$ which is due to Drude resistance only, i.e., what is expected in the absence of B-field and hydrodynamics ($\dot{E}_{0}/LW=nm^{*}v_{0}^{2}/\tau$). The value $\omega\tau=1.2$ corresponds to $B_{0}=75$ mT and $\omega\tau=2.4$ corresponds to $B_{0}=150$ mT. In these plots we assume that the viscosity is independent of the magnetic field, $\nu(B)=\nu(0)$. The plots correspond to the parameter set (II), where $gD=1.16$.
However, the viscosity depends on magnetic field, with possible dependence [10]
$\nu\propto\frac{1}{1+(B/B^{*})^{2}}\;,\;B^{*}=\frac{p_{F}}{2|e|l_{ee}}$ (22)
For the device parameters (II) the characteristic magnetic field is $B^{*}\approx 240\>\text{mT}$. Therefore, for $B_{0}=150$ mT the approximation $\nu(B)=\nu(0)$ is questionable, but the possible correction does not exceed 10-20%. The viscous dissipation comes from the regions of maximum vorticity and in this setup the vorticity is maximum at the regions where $B\to 0$, so the true correction for $B_{0}=150$ mT is lower than that given simply by Eq. (22). It is not a problem to solve the Navier-Stokes equation with position dependent viscosity, but it is not necessary at this stage.
@@
Interestingly, the fluid flow in the Knudsen layer shows a sort of anti-Poiseuille behaviour. To illustrate this, we plot in Fig. 3 the longitudinal velocity $v_{x}$ versus $y$ (i.e., across the channel). The velocity $v_{x}$ is a function of both $x$ and $y$ and here we present plots at $x=0$. At a sufficiently large magnetic field, $\omega\tau\geq 2.4$, there is a clear anti-Poiseuille profile near the left boundary of the channel. In this regime the peak velocity is not $v_{0}$ in the bulk, but instead the longitudinal velocity has a maximum in the boundary layer before decaying towards $v_{0}$ in the bulk. There is a conventional Poiseuille behaviour at the right boundary of the channel, although at high magnetic field a small amount of counterflow develops due to the development of vortices in the boundary layer. If we shift $x=0\to x=a/2$ the plots are effectively reflected about the centre of the channel: the anti-Poiseuille profile is located at the right boundary and the Poiseuille profile is at the left boundary. Mathematically, the anti-Poiseuille profile is related to the complex eigenvalues $\lambda_{a}$. This behaviour is qualitatively similar to the anti-Poiseuille flow predicted for narrow graphene channels near charge neutrality in the presence of a uniform magnetic field perpendicular to the plane [49, 50, 51]. Near charge neutrality this effect is largely attributed to electron-hole recombination. In our case the physical origin of the anti-Poiseuille profile is different as we do not have any holes. While we consider no-slip boundaries, a similar anti-Poiseuille profile exists also for boundary conditions with a finite slip length. The location of the anti-Poiseuille maximum may be used to extract the slip length in a sample with finite-slip boundaries.
@@
Figure 3: Longitudinal velocity profile at $x/D=0$ with $gD=1.16$ for various values of $\omega\tau$. At large magnetic fields anti-Poiseuille flow is observed and a small counterflow is observed at the opposite boundary due to the formation of a viscous vortex.
@@ V Finite-elements simulations
Figure 4: FEM results for the absolute value $|\mathbf{v}|$ of the velocity in the channel for $gD=0.5$. Note that this color plot should not be confused with lines of flow, which are shown in Fig. 1 The velocity is measured in units convenient for numerical solution and the system contains six magnetic strips. The region of high velocity, anti-Poiseuille flow, is shifted by half a period at opposite boundaries.
@@ VI Magnetic field of periodic ferromagnetic stripes
Figure 5: Left: top view of a periodic array of magnetic bars perpendicular to the flow direction shown by the red arrow. The period of the array is $a$ and the width of the bar is $d$. The bars are magnetized in parallel to the longitudinal axis of the channel. The direction of the magnetization is shown by the blue arrows. Right: Side view of the device. The vertical distance from the bar magnets to the 2DEG is $z$ and the thickness of the rectangular bar is $\delta z$.
The bars are magnetized along the longitudinal axis of the channel. The bar thickness is $\delta z$ and the distance from the base of the bar to the 2DEG is $z$, shown in the right panel of Fig. 5. The period of modulation is $a=1.9$ $\mu$m as found in Sec. II. We tune the remaining free parameters $d,z,\delta z$ such that the $z$ component of the magnetic field in the 2DEG is approximately $B_{0}\sin(2\pi x/a)$ with $B_{0}\approx 75$ mT. A stronger magnetic field may be also achieved.
Assuming the magnetization is uniform throughout the volume of each bar, and the bars’ extent in the transverse ($y$) direction is wider than the width $W$ of the sample, we calculate the magnetic field in the plane of the 2DEG as a function of longitudinal coordinate $x$ from the Biot-Savart law as
@@
where $B_{m}$ is the saturation field, and $x$ is measured from the center of one of the magnets. For example, at the edge of the magnet $x=d/2$, the harmonics of the field in the 2DEG plane is $\propto\sin^{2}(gnd/2)$. Therefore, $d=a/2=\pi/g$ maximizes the first harmonic and suppresses the second harmonic. We numerically calculate $B$ from Eq. (VI) and further confirm that $d=a/2=0.93\>\mu$m is optimal.
In order to achieve a field with magnitude $B_{0}=75$ mT we consider a NiFe alloy that has saturation magnetic field $1500$ mT. [56] Through concurrently tuning $\delta z$ and $z$ we find that we may generate the desired magnetic field profile with $\delta z=0.15$ $\mu$m and $z=0.30$ $\mu$m. Plots of $B_{z}$ versus $x$ and $B_{x}$ versus $x$ are presented in Fig. 6.
Figure 6: Magnetic field in the 2DEG plane for NiFe bars. Left panel: $B_{z}$ versus $x$. Right panel: $B_{x}$ versus $x$. The parameters are $\delta z=0.15\>\mu\mathrm{m},\ z=0.30\>\mu\mathrm{m}$.
The $x$ component of the magnetic field is not relevant, but we present it for completeness. The $x$ dependence of $B_{z}$ is not quite sinusoidal, however this is not a big issue as using the finite element method one can also solve the Navier-Stokes equation for this magnetic field profile.
@@ VII Conclusions
We develop a theory for the no-slip boundary layer in the magnetohydrodynamic flow of an electron fluid in a straight channel. The fluid flow is modified by an array of micromagnets on the top of two-dimensional electron gas. The micromagnets create a wiggling flow of the fluid, which dramatically enhances hydrodynamic effects in the fluid. To be specific, the analysis was performed for no-slip boundary conditions, but the method can be extended to boundary conditions with arbitrary slip length. We demonstrate that the fluid velocity profile in the boundary layer manifests an anti-Poiseuille behaviour. In addition, we show that the micromagnet array creates a longitudinal Hall effect in the sample. From the experimental point of view we propose a method for a precise measurements of the properties of different electron fluids. The results are applicable to graphene away from the charge neutrality point and also to semiconductors.
@@ Appendix A The method for exact solution for a finite width channel
$m(f_{m}e^{\mp W/2}+h_{m}e^{\pm W/2}+\sum_{a}(\alpha_{a}A_{m,a}^{(n)})e^{\mp\lambda_{a}W/2}$
$+\gamma_{a}C_{m,a}^{(n)}e^{\pm\lambda_{a}W/2}))+\frac{\omega\tau}{2g(1+g^{2})}(\delta_{m,1}-\delta_{m,-1})=0$
@@
$\sum_{a}\lambda_{a}(\alpha_{a}A_{m,a}^{(n)}e^{\mp\lambda_{a}W/2}-\gamma_{a}C_{m,a}^{(n)})$
$+f_{m}e^{\mp W/2}-h_{m}e^{\pm W/2}+e^{\pm\lambda_{a}W/2}=0$ (27)

### v2

@@ III Boundary layer for a straight semi-infinite channel with no-slip boundary condition
This section has two goals, (i) to estimate the width of the boundary layer, (ii) to point out some qualitative effects related to the boundary layer. In this section we disregard the subleading gradient correction and use the standard Navier-Stokes equation (II).
@@ III.1 Boundary layer flow in the perturbative limit $\omega\tau\ll 1$
The no-penetration condition ensures that no fluid can cross the boundary whereas the no-slip condition ensures that the fluid has zero tangential velocity at the boundary surface. Furthermore, the stream function must decay to the bulk solution (II) far from the boundary. We look for a solution as an expansion in powers of $\omega\tau\ll 1$, i.e., $\psi=\sum_{n=0}^{\infty}(\omega\tau)^{n}\psi^{(n)}$. First we solve for the zero-order solution, corresponding to flow in the absence of a magnetic field. Solving the Stokes equation for $\omega\tau=0$ using the boundary conditions leads to the Poiseuille solution
@@ III.2 Exact numerical solution for the boundary layer flow
Refer to captionRefer to captionRefer to caption
Figure 1: Streamlines in a semi-infinite channel with a no-slip boundary at $y=0$. The parameters are chosen as $gD=0.88$, $\omega\tau=0.2,1.2,2.4$. Solid lines correspond to the numerically exact solution described in Sec. III.2 and dashed lines correspond to the solution given by Eq. (II) in the infinite device.
@@
where $f_{m}$ and $A_{m,a}$ are complex coefficients. The index $a=1,2,3...$ enumerates “eigenvalues” $\lambda_{a}$, which can be real or complex. The real parts of $\lambda_{a}$ are positive by design of Eq. (17). In practice the upper limit of the index $a$, $a_{max}=2N_{m}+1$, is fixed by truncation. Since the “Bernoulli term” is neglected, the Stokes equation (II) is a linear equation. Let us consider the Poiseuille solution without magnetic field,
@@
The boundary layer solution is extended to finite width channels in App. A. After the method is developed it is instructive to present plots for specific cases. For the numerical solution we set $gD=0.88$, which corresponds to the device described at the end of Sec. II. We truncate beyond $N_{m}=5$, which provides more than sufficient accuracy for the values of $\omega\tau$ we present. Plots of the streamlines obtained for $\omega\tau=0.2$, $1.2$ and $2.4$ are presented in Fig. 1. These values of $\omega\tau$ were selected as $\omega\tau=0.2$ has only real eigenvalues $\lambda_{a}$ and corresponds to the perturbative regime, $\omega\tau=1.2$ has a complex conjugate pair of eigenvalues $\lambda_{a}$ and corresponds to the device described in Eq. (II) and $\omega\tau=2.4$ has two complex conjugate pairs of eigenvalues $\lambda_{a}$. As the magnetic field increases ($\omega\tau$ increases) both the magnitude of the boundary effects and the width of the boundary layer increase. For the most interesting case $\omega\tau\approx 1$ the effective width of the boundary layer is approximately $D$. Therefore, for the set of parameters (II) a channel of width $=10-20\mu$m is suffiently wide to neglect boundary layer effects.
The boundary layer solution considered in this Section is valid for the no-slip boundary condition. However, the method can be easily extended to the case of an arbitrary slip length, see App. B for details.
@@ IV Numerical estimates, anti-Poiseuille flow
Following the numerical calculation of the stream function, we may calculate the the viscous dissipation rate and the Ohmic dissipation rate given by Eqs. (II). The most important parameter is the ratio $\dot{E}_{\nu}/\dot{E}_{\Omega}$, which is dependent on the device considered, so to provide calculated values we take the device with parameters presented in Eq. (II). This corresponds to $D=0.34\>\mu\text{m}$. We also take the lattice spacing $a=2.5\>\mu\text{m}$ as discussed in the end of Section II, resulting in $gD=0.88$. The dissipation rates are plotted in Fig. 2 versus the channel width $W$ for two values of the magnetic field amplitude, $B_{0}=75$ mT, $150$ mT, corresponding to $\omega\tau=1.2$, $2.4$ respectively. Solid lines correspond to solutions of the NS equation (II). The dashed line corresponds to total dissipation in the bulk from solution of the NS+ equation (3) where the gradient correction is significant, $C^{2}=1\to C^{2}=(1-0.19)^{2}=1-0.34$. In these plots we assume that the viscosity is independent of the magnetic field, $\nu(B)=\nu(0)$. Asymptotic values of the ratio $\dot{E}_{\nu}/\dot{E}_{\Omega}$ at $W\to\infty$ are $0.16$ ($B_{0}=75$ mT) and $0.48$ ($B_{0}=150$ mT). Note that as magnetic field increases the difference between the asymptotic value of the total dissipation from the NS equation and the total bulk dissipation from the NS+ equation increases.
Refer to captionRefer to caption
@@
Figure 2: Ohmic, viscous and total dissipation from solution of the NS equation (II) as a function of channel width and total dissipation in the bulk from solution of the NS+ equation (3). The dissipation is normalized to the $\dot{E}_{0}$ which is due to Drude resistance only, i.e., what is expected in the absence of B-field and hydrodynamics ($\dot{E}_{0}/LW=nm^{*}v_{0}^{2}/\tau$). The value $\omega\tau=1.2$ corresponds to $B_{0}=75$ mT and $\omega\tau=2.4$ corresponds to $B_{0}=150$ mT. In these plots we assume that the viscosity is independent of the magnetic field, $\nu(B)=\nu(0)$. The plots correspond to the parameter set (II), where $gD=0.88$.
Now let us calculate the value of the correction of the first kind as it is discussed in the paragraph before Eq.(3). The viscosity is often reported to depend on magnetic field, with possible dependence [10]
$\nu\to\frac{\nu}{1+(B/B^{*})^{2}}$ $\approx\nu[1-(B/B^{*})^{2}]\;,\;$
$B^{*}$ $=\frac{p_{F}}{2|e|l_{ee}}$ (25)
The term $\nu(B/B^{*})^{2}\nabla^{2}\mathbf{v}$ added to the Navier-Stokes Eq. (II) leads to the correction of the first kind. we call it ”the amplitude correction”. It is easy calculate the correction by perturbation theory. In perturbation theory it does not influence the Ohmic dissipation, but in calculating the viscous dissipation in Eq.(II), as $B=B(x)$ gives $\nu=\nu(x)$ one has to move $\nu$ in the integrand and perform integration over position. This gives the following correction to the viscous dissipation.
$\frac{\delta{\dot{E}}_{\nu}}{{\dot{E}}_{\nu}}=-3(\omega\tau)^{2}\left(\frac{l_{ee}}{l_{mfp}}\right)^{2}$ (26)
The correction is $-7\%$ for $\omega\tau=1.2$ and $-28\%$ for $\omega\tau=2.4$. Note that by performing experiments for different values of $a$ and different values of the magnetic field amplitude it is possible to disentangle and to measure independently the viscosity, the gradient correction (Hall viscosity), and the amplitude correction (26).
@@
Interestingly, the fluid flow in the boundary layer shows a sort of anti-Poiseuille behaviour. To illustrate this, we plot in Fig. 3 the longitudinal velocity $v_{x}$ versus $y$ (i.e., across the channel). The velocity $v_{x}$ is a function of both $x$ and $y$ and here we present plots at $x=0$. At a sufficiently large magnetic field, $\omega\tau\geq 1.2$, there is a clear anti-Poiseuille profile near the left boundary of the channel. In this regime the peak velocity is not $v_{0}$ in the bulk, but instead the longitudinal velocity has a maximum in the boundary layer before decaying towards $v_{0}$ in the bulk. There is a conventional Poiseuille behaviour at the right boundary of the channel, although at high magnetic field a small amount of counterflow develops due to the development of vortices in the boundary layer. If we shift $x=0\to x=a/2$ the plots are effectively reflected about the centre of the channel: the anti-Poiseuille profile is located at the right boundary and the Poiseuille profile is at the left boundary. Mathematically, the anti-Poiseuille profile is related to the complex eigenvalues $\lambda_{a}$. This behaviour is qualitatively similar to the anti-Poiseuille flow predicted for narrow graphene channels near charge neutrality in the presence of a uniform magnetic field perpendicular to the plane [50, 51, 52]. Near charge neutrality this effect is largely attributed to electron-hole recombination. In our case the physical origin of the anti-Poiseuille profile is different as we do not have any holes. While we consider no-slip boundaries, a similar anti-Poiseuille profile exists also for boundary conditions with a finite slip length, see Appendix B.
@@
Figure 3: Longitudinal velocity profile at $x/D=0$ with $gD=0.88$ for various values of $\omega\tau$. At large magnetic fields anti-Poiseuille flow is observed and a small counterflow is observed at the opposite boundary due to the formation of a viscous vortex.
@@ V Finite-elements simulations
Figure 4: FEM results for the absolute value $|\mathbf{v}|$ of the velocity in a finite channel of length $L=100D$ for $gD=0.5$. There is an equilibration region near the inlet and outlet where uniform flow transitions to periodic flow, with periodic flow established from $-40<x/D<40$. Note that this color plot should not be confused with lines of flow, which are shown in Fig. 1 The velocity is measured in units convenient for numerical solution and the system contains six magnetic strips. The region of high velocity, anti-Poiseuille flow, is shifted by half a period at opposite boundaries.
@@ VI Magnetic field of periodic ferromagnetic stripes
Figure 5: Left: top view of a periodic array of magnetic bars perpendicular to the flow direction shown by the red arrow. The micromagnets are magnetized parallel to the longitudinal axis of the channel. The direction of the magnetization is shown by the blue arrows. The period of the superlattice is $a$ and the width of each micromagnet is $d$. Right: Side view of the device. The vertical distance from the micromagnets to the 2DEG is $z$ and the thickness of each micromagnet is $\delta z$. First presented in Ref. [53].
The thickness of each micromagnet is $\delta z$ and the distance from the base of each micromagnet to the 2DEG is $z$, shown in the right panel of Fig. 5. The period of modulation is $a=2.5$ $\mu$m as found in Sec. II. The micromagnets are magnetized parallel to the longitudinal axis of the channel. The remaining free parameters $d,z,\delta z$ are tuned such that the $z$ component of the magnetic field in the plane of the 2DEG is approximately $B_{0}\sin(2\pi x/a)$ with $B_{0}\approx 75$ mT. A stronger magnetic field may be also achieved.
Assuming the magnetization is uniform throughout the volume of each micromagnet, and each micromagnet’s extent in the transverse ($y$) direction is wider than the width $W$ of the sample, we calculate the $z$ component of the magnetic field in the plane of the 2DEG as a function of longitudinal coordinate $x$ from the Biot-Savart law as
@@
where $B_{m}$ is the saturation magnetic field, and $x$ is measured from the center of one of the micromagnets. At the edge of the micromagnet centred on $x=d/2$, the harmonics of the field in the 2DEG plane is $\propto\sin^{2}(gnd/2)$, so evidently $d=a/2=\pi/g$ maximizes the first harmonic and suppresses the second harmonic. We calculate $B$ numerically from Eq. (VI) and confirm that $d=a/2=1.25\>\mu$m is optimal.
In order to achieve a field with magnitude $B_{0}=75$ mT we consider a NiFe alloy that has saturation magnetic field $1500$ mT. [58] Through concurrently tuning $\delta z$ and $z$ we find that we may generate the desired magnetic field profile with $\delta z=0.20$ $\mu$m and $z=0.40$ $\mu$m. Plots of $B_{z}$ versus $x$ and $B_{x}$ versus $x$ are presented in Fig. 6.
Figure 6: Magnetic field in the 2DEG plane for NiFe bars. Left panel: $B_{z}$ versus $x$. Right panel: $B_{x}$ versus $x$. The parameters are $\delta z=0.20\>\mu\mathrm{m},\ z=0.40\>\mu\mathrm{m}$.
The $x$ component of the magnetic field is not relevant, but we present it for completeness. The $x$ dependence of $B_{z}$ is not quite sinusoidal, however this is not a big issue as using the finite element method one can also solve the Navier-Stokes equation for this magnetic field profile. It should be noted that any magnetic disorder due to a slight misalignment of the micromagnets with respect to the channel or inconsistencies in the spacing or width of the micromagnets results in high wavenumber contributions to the magnetic field, with these contributions exponentially suppressed with due to the distance $z$ between the micromagnets and the 2DEG as seen in Eq. (VI).
@@ VII Conclusions
We develop a theory for the no-slip boundary layer in the magnetohydrodynamic flow of an electron fluid in a straight channel. The fluid flow is modified by an array of micromagnets on the top of two-dimensional electron gas. The micromagnets create a wiggling flow of the fluid, which dramatically enhances hydrodynamic effects in the fluid. To be specific, the analysis was performed for no-slip boundary conditions, but the method can be extended to boundary conditions with arbitrary slip length. We demonstrate that the fluid velocity profile in the boundary layer manifests an anti-Poiseuille behaviour. In addition, we show that the micromagnet array creates a longitudinal voltage modulation in the sample. From the experimental point of view we propose a method for a boundary-independent measurement of the viscosity of different electron fluids. The results are applicable to graphene away from the charge neutrality point and also to semiconductors.
@@ Appendix A The method for exact solution for a finite width channel
$m\bigg{[}f_{m}e^{\mp W/2}+h_{m}e^{\pm W/2}+\sum_{a}(\alpha_{a}A_{m,a}^{(n)}e^{\mp\lambda_{a}W/2}$
$+\gamma_{a}C_{m,a}^{(n)}e^{\pm\lambda_{a}W/2})\bigg{]}+\frac{\omega\tau}{2g(1+g^{2})}(\delta_{m,1}-\delta_{m,-1})=0$
@@
$\sum_{a}\lambda_{a}(\alpha_{a}A_{m,a}^{(n)}e^{\mp\lambda_{a}W/2}-\gamma_{a}C_{m,a}^{(n)}e^{\pm\lambda_{a}W/2})$
$+f_{m}e^{\mp W/2}-h_{m}e^{\pm W/2}=0$ (31)
@@
Appendix B Solution for Boundaries With Finite Slip Length
We now expand upon the solution for the finite channel to account for some finite slip length, $\beta$, defined in the finite-slip boundary condition.[59] Note the change in sign of $\beta$ at the positive and negative $y$ boundaries.
$v_{x}|_{y=\pm W/2}=\mp\beta\partial_{y}v_{x}|_{y=\pm W/2}$ (32)
The no-penetration boundary condition remains unchanged. We must first account for the finite-slip boundary in our definition of $\psi_{0}$, which we do by considering flow in the absence of perioidic magnetic field. This is simply the following.
$\psi_{0}=y+\frac{e^{-y}-e^{y}}{e^{w/2}(1+\beta)+e^{-w/2}(1-\beta)}$ (33)
While the equations for the free solution, Eq. (A) remain unchanged, the equations for the driven solution, Eq. (29) are modified by the redefinition of $\psi_{0}$.
$(g^{2}m^{2})(g^{2}m^{2}-1)f_{m}+\frac{\omega\tau g}{2}(f_{m+1}+f_{m-1})$ (34)
$=-\frac{\omega\tau g}{2(e^{w/2}(1+\beta)+e^{-w/2}(1-\beta))}(\delta_{m,1}+\delta_{m,-1})$
$(g^{2}m^{2})(g^{2}m^{2}-1)h_{m}-\frac{\omega\tau g}{2}(h_{m+1}+h_{m-1})$
$=-\frac{\omega\tau g}{2(e^{w/2}(1+\beta)+e^{-w/2}(1-\beta))}(\delta_{m,1}+\delta_{m,-1})$
Now, while the no-penetration boundary condition remains unchanged, we must solve the finite-slip boundary condition for $\delta\psi_{d}+\delta\psi_{f}$. The no-penetration and finite slip boundary conditions applied to the finite channel system are as follows.
$0=v_{y}|_{y=\pm W/2}\to$
$m\bigg{[}f_{m}e^{\mp W/2}+h_{m}e^{\pm W/2}+\sum_{a}(\alpha_{a}A_{m,a}^{(n)}e^{\mp\lambda_{a}W/2}$
$+\gamma_{a}C_{m,a}^{(n)}e^{\pm\lambda_{a}W/2})\bigg{]}+\frac{\omega\tau}{2g(1+g^{2})}(\delta_{m,1}-\delta_{m,-1})=0$
$v_{x}|_{y=\pm W/2}=\mp\beta\partial_{y}v_{x}|_{y=\pm W/2}\to$
$\sum_{a}\lambda_{a}(\alpha_{a}A_{m,a}^{(n)}e^{\mp\lambda_{a}W/2}-\gamma_{a}C_{m,a}^{(n)}e^{\pm\lambda_{a}W/2})$
$+f_{m}e^{\mp W/2}-h_{m}e^{\pm W/2}=\mp\beta\bigg{[}f_{m}e^{\mp W/2}+h_{m}e^{\pm W/2}$
$+\sum_{a}\lambda_{a}^{2}(\alpha_{a}A_{m,a}^{(n)}e^{\mp\lambda_{a}W/2}+\gamma_{a}C_{m,a}^{(n)}e^{\pm\lambda_{a}W/2})\bigg{]}$ (35)
Figure 7: Longitudinal velocity profile at $x/D=0$ with $gD=0.88$ and $\omega\tau=4.8$ for a range of slip length $\beta$. Clearly the finite-slip cases interpolate between the no-slip and perfect-slip cases. As $\beta$ increases the strength of the anti-Poiseuille flow peak increases until it saturates at the perfect-slip level, while counterflow at the other boundary quickly disappears with finite $\beta$.
It is instructive to compare the no-slip, finite-slip and perfect-slip cases. For monolayer graphene the slip length is dependent on temperature and has been found, as a fitting parameter, to be on the order of $0.5-0.1$ $\mu$ m for $T=300$K[22] and $T=75K$[23]. For the parameters we consider in this manuscript an appropriate choice is $\beta=D/3$. In Fig. 7 we present $v_{x}$ profiles for $\omega\tau=4.8$ for the different boundary conditions, with this $\omega\tau$ chosen to highlight anomalous flow behaviour. As the $\beta$ is increased the flow transitions from the no-slip case to the perfect-slip case, with the anti-Poiseuille flow peak increasing with increasing $\beta$ and counterflow at the opposite boundary quickly vanishing with small finite $\beta$. At realistic slip length $\beta=D/3$ one would expect strong periodic anti-Poiseuille flow as well as small finite counterflow.